# MobiFuse: Learning Universal Human Mobility Patterns through Cross-domain Data Fusion

### v1

@@ Abstract
Human mobility modeling is critical for urban planning and transportation management, yet existing datasets often lack the resolution and semantic richness required for comprehensive analysis. To address this, we proposed a cross-domain data fusion framework that integrates multi-modal data of distinct nature and spatio-temporal resolution, including geographical, mobility, socio-demographic, and traffic information, to construct a privacy-preserving and semantically enriched human travel trajectory dataset. This framework is demonstrated through two case studies in Los Angeles (LA) and Egypt, where a domain adaptation algorithm ensures its transferability across diverse urban contexts. Quantitative evaluation shows that the generated synthetic dataset accurately reproduces mobility patterns observed in empirical data. Moreover, large-scale traffic simulations for the LA County based on the generated synthetic demand align well with observed traffic. On California’s I-405 corridor, the simulation yields a Mean Absolute Percentage Error of 5.85% for traffic volume and 4.36% for speed compared to Caltrans PeMS observations.
Data Fusion, Domain Adaptation, Human Mobility Pattern, Travel Trajectory
@@ 1.1 Motivation
Human mobility modeling has emerged as a critical component in urban planning, management of transportation, business, and public policy development (González et al., 2008; Liao et al., 2024a; Stanford et al., 2024). Understanding how people move through cities and interact with the surrounding environment is fundamental to creating efficient, sustainable, and livable urban spaces. This importance has grown with increasing urbanization and the complexities of modern transportation systems (Hasan et al., 2013; Ma et al., 2024b).
One of the key research directions in mobility modeling lies in integrating heterogeneous data sources to address incomplete or sparse datasets. Data fusion has emerged as a critical approach for combining diverse data types and enhancing modeling accuracy and robustness. Existing methods have demonstrated the potential of integrating sources such as GPS, social media, and environmental data to provide richer insights into mobility patterns (Ounoughi & Yahia, 2023; Afyouni et al., 2022). Researchers have also explored techniques for cross-domain integration and spatial-temporal modeling to improve movement predictions and scalability (Zheng, 2015; Wei et al., 2024). Although significant progress has been made, existing approaches still encounter challenges such as data sparsity, heterogeneity, and limited adaptability across diverse geographic contexts. To address these issues, this paper proposes a context-aware data fusion framework that enhances the applicability, precision, and generalizability of mobility modeling across diverse geographic and socio-demographic settings.
Figure 1: The need for data fusion and the fused data.
@@ 1.2 Challenges
Comprehensive datasets addressing the diverse requirements of mobility research remain scarce. An ideal mobility dataset would integrate rich semantic information with high-resolution trajectory data and demographic attributes while ensuring privacy protection. However, as shown in Figure 1 (a), existing datasets often focus on only a subset of these dimensions (Kapp & Mihaljevic, 2023). The complexity of combining these diverse data sources stems from two key challenges: cross-domain data harmonization and multi-modal information integration, as shown in Figure 2.
The fundamental challenge of cross-domain data fusion lies in combining data sources that complement each other but often contain conflicting information. Our framework addresses this by fusing data listed in Figure 2, where each source has distinct limitations: trajectories lack semantic context and are usually low quality, surveys are biased and expensive, and infrastructure data requires synthesis with human mobility patterns to be meaningful.
@@
To address these challenges, we propose a cross-domain data fusion framework that synthesizes mobility data with integrated semantic, trajectory, demographic, and location-based information, as shown in Figure 1 (b). By leveraging the complementary strengths of these diverse data sources, our approach enables a more comprehensive understanding of mobility patterns while preserving privacy.
@@ 1.3 Contribution
Compared to prior work in data fusion and mobility modeling, we bridge multiple data gaps between: high-resolution but semantically-sparse trajectories and context-rich but limited surveys, population-level statistics and individual-level behaviors, and static infrastructure data and dynamic human movements, enabling new insights into human mobility and its implications for societal infrastructure, which hold promise for numerous applications, including traffic optimization, urban planning, and anomaly detection.(He et al., 2024; Ma et al., 2024a; Stanford et al., 2024).
Compared with the existing literature, our research makes several significant contributions:
  • We proposed a novel data fusion framework that processes and integrates cross-domain multi-modal data sources to create comprehensive synthetic datasets, while balancing usability and privacy considerations. The framework harmonizes heterogeneous data types (spatial-temporal data: GPS (Veraset, 2024) and Point of Interest (POI) (OpenStreetMap contributors, 2017), mobility data: Household Travel Survey (HTS) (Administration, 2017), demographic data: Time Use Survey (TUS) (U.S. BLS, 2024), and traffic data: road network data (OpenStreetMap contributors, 2017)) and finally validated with large-scale traffic simulation.
  • We developed a methodology that fuses the realism of GPS data with the knowledge of survey data by using semantic annotations from an LLM and reconstructing missing activities with a survey-trained model, improving the completeness and interpretability of mobility trajectories.
  • We proposed a transfer learning framework that addresses the two key data challenges: 1) enabling knowledge and pattern adaptation in data-sparse regions through semi-supervised learning, leveraging limited GPS trajectories to adapt source domain patterns without requiring complete ground truth data, and 2) incorporating demographic variations into anonymized trajectories through unsupervised distribution adaptation, aligning mobility patterns with demographic characteristics without pairwise individual labels.
Figure 2: An overview of the proposed data fusion framework. The captions within methodology indicate their corresponding sections in the paper for reference.
@@ 2.1 Human Travel Trajectory Reconstruction
Trajectory reconstruction has become crucial in understanding human mobility patterns, especially when dealing with incomplete datasets. Recent advancements have addressed the challenges of data sparsity and irregularity through innovative techniques. Chen et al. (Chen et al., 2019) introduced the Context-enhanced Trajectory Reconstruction (CTR) method, using tensor factorization to reconstruct complete individual trajectories from sparse Call Detail Records. Li et al. (Li et al., 2019) proposed the Multi-criteria Data Partitioning Trajectory Reconstruction (MDP-TR) method for large-scale, low-frequency mobile phone datasets, enhancing reconstruction performance by considering spatiotemporal patterns of missing data and individual similarities. For GPS data, Zheng et al. (Zheng et al., 2010) developed a collaborative system for location and activity recommendations, demonstrating significant improvements in inferring activity types. Alexander et al. (Alexander et al., 2015) emphasized the importance of comprehensive temporal and spatial analysis in trajectory reconstruction using mobile phone data.
@@ 2.2 Large Language Models for Mobility Modeling
Large language models (LLMs) have demonstrated significant potential in mobility modeling by leveraging their ability to process contextual and semantic information from diverse data sources. Recent advancements, such as those by Liu et al. (Liu et al., 2024b) and Cai et al. (Cai et al., 2024), highlight the versatility of LLMs in tasks like trajectory reconstruction and decision-making. Studies on foundational models like LLaMA (Touvron et al., 2023), PaLM (Anil et al., 2023), and TEG-DB (Li et al., 2024) further underline their power in multi-modal data fusion and semantic interpretation. These capabilities make LLMs valuable complements to traditional methods, particularly in addressing challenges of data sparsity and complex semantic understanding in mobility applications.
@@ 2.3 Transfer Learning
Howard and Ruder’s (Howard & Ruder, 2018) gradual unfreezing technique could be adapted to preserve general mobility patterns while adjusting to specific regional characteristics. Peters et al. (Peters et al., 2019) and Merchant et al. (Merchant et al., 2020) suggest that middle layers are often the most transferable, which could be useful when adapting mobility models from data-rich to data-scarce regions. Liu et al. (Liu et al., 2019) explored freezing bottom layers while fine-tuning top layers, a strategy that could preserve fundamental human movement patterns while adapting to unique characteristics of specific urban environments.
@@ 3 Problem Formulation
The transfer learning stage (4.2) is designed to capture real-world mobility dynamics while enabling deployment in data-scarce regions. Through knowledge (4.2.1) and pattern (4.2.2) adaptation, the model can be adapted to different regions and cultures and learns to balance the generic knowledge learned from surveys with the realism in GPS trajectories, including region-specific patterns. Our demographic adaptation module further enhances realism by incorporating sociodemographic variations using TUS data, which provides rich activity patterns across different population groups, ensuring our synthetic trajectories reflect both demographic-specific behaviors and real-world mobility dynamics while maintaining privacy.
Finally, the third stage (4.3) grounds these behaviorally-realistic patterns in physical space by fusing the generated activity chains with spatial distributions and traffic data to assign feasible locations, enabling validation through large-scale traffic simulation. This complete integration of dynamic human behavior with static infrastructure data creates a comprehensive framework that generates mobility patterns reflecting both the complexity of human decision-making and the constraints of physical infrastructure.
@@
With our semantically annotated trajectories represented as $A_{j}$, we propose an activity reconstruction model $M^{R1}$ tailored to a specific region $R1$. For an incomplete $A_{j}$, the model synthesis a new activity chain $A_{j}^{\prime}=M^{R1}(A_{j})$, by filling the gaps based on the common activity patterns learned from the region $R1$.
The activity reconstruction adapts cultural effects to activity patterns through transfer learning, but incorporating sociodemographic correlation patterns poses a challenge. To address this, we introduce a demographic adaptation module that modifies reconstructed activity chains based on sociodemographic characteristics. Let us define the sociodemographic space $\mathcal{S}$ as: $\mathcal{S}={\mathbf{s}_{g}|g\in\mathcal{G}}$ where $\mathcal{G}$ represents the set of demographic groups. For a demographic space with $N$ dimensions (e.g., age, gender, employment status), where each dimension $i$ has $C_{i}$ categories, the total number of demographic groups is: $|\mathcal{G}|=\prod_{i=1}^{N}C_{i}$.
The adapter functions as an additional layer after the reconstruction model $M^{R1}$. Given a batch of reconstructed activity chains probability matrix $\mathbf{Y}\in\mathbb{R}^{B\times K\times|T|}$, where $B$ is the batch size, $K=96$ represents 96 time steps with 15-minute time intervals for a day, and $|T|$ represents the number of activity types defined in Table 1, we introduce an adapter matrix $\mathbf{M}\in\mathbb{R}^{|\mathcal{G}|\times|T|}$. For each demographic group $g\in\mathcal{G}$, its adapter vector $\mathbf{a}_{g}\in\mathbb{R}^{|T|}$ modifies the reconstructed activity probabilities. The demographically-adapted activity probabilities for group $g$ across $K$ time intervals are computed as: $\mathbf{P}_{g}=\text{softmax}(\mathbf{Y}+\mathbf{a}_{g},\text{axis}=|T|)$, where $\mathbf{a}_{g}$ is broadcast across the $K$ time intervals.
Finally, we assign the activity chains to physical zonal locations while preserving spatial and temporal feasibility. Given a set of TAZs $\mathcal{Z}$ and road network $\mathcal{R}$, we assign locations to activities by optimizing:$L^{*}=\operatorname*{arg\,min}_{L\in\mathcal{Z}}\sum_{i=1}^{n}\sum_{j=1}^{m_{i}}d(L_{i,j},L_{i,j+1})$, subject to constraints on travel time feasibility via $\mathcal{R}$ and consistency with observed spatial distributions. This enables validation against real-world traffic patterns.
@@ 4.1 Generative Data Fusion
Human trajectory data collected from GPS devices include geo-location and time-stamped stay points that elucidate individual mobility patterns. However, these trajectories suffer from both semantic sparsity and temporal incompleteness. We develop the generation based data fusion module, as Figure 2 (4.1), to address these two challenges.
Adopting the methodology outlined (Liu et al., 2024a), we use LLMs to annotate stay points using the POI dataset. This process utilizes LLM to assign semantic labels to POIs based on attributes such as ”name”, ”amenity”, and ”construction type”. Each POI is then categorized into one of 15 activity codes as described in Table 1. The LLM’s prompt configuration includes a ”Task Description”, ”Activity Code Illustration”, and a description of each POI in natural language format, ensuring comprehensive understanding and accurate classification. The output from the LLM provides the top three most relevant activity codes for each POI, each with a corresponding probability.
@@
While this LLM-based annotation provides semantic context to the trajectories, their inherent incompleteness limits their direct use for mobility analysis. Therefore, our framework continues with an activity reconstruction process that fuses these annotated but incomplete trajectories with comprehensive mobility data from travel surveys. A base Activity Reconstruction Model $M^{0}$ is trained on survey data to learn generic activity patterns and can be adapted to other regions with small dataset, supported by the strong similarity in human activities across regions  (Schneider et al., 2013; Cao et al., 2019). The adaptation of this base model to realistic geographical and cultural contexts is detailed in Section 4.2.

### v2

# Conceptualization of this study, Methodology, Implementation, Writing

@@ Abstract
Human mobility modeling is critical for urban planning and transportation management, yet existing approaches often lack the integration capabilities needed to handle diverse data sources. We present a foundation model framework for universal human mobility patterns that leverages cross-domain data fusion and large language models to address these limitations. Our approach integrates multi-modal data of distinct nature and spatio-temporal resolution, including geographical, mobility, socio-demographic, and traffic information, to construct a privacy-preserving and semantically enriched human travel trajectory dataset. Our framework demonstrates adaptability through domain transfer techniques that ensure transferability across diverse urban contexts, as evidenced in case studies of Los Angeles (LA) and Egypt. The framework employs LLMs for semantic enrichment of trajectory data, enabling comprehensive understanding of mobility patterns. Quantitative evaluation shows that our generated synthetic dataset accurately reproduces mobility patterns observed in empirical data. The practical utility of this foundation model approach is demonstrated through large-scale traffic simulations for LA County, where results align well with observed traffic data. On California’s I-405 corridor, the simulation yields a Mean Absolute Percentage Error of 5.85% for traffic volume and 4.36% for speed compared to Caltrans PeMS observations, illustrating the framework’s potential for intelligent transportation systems and urban mobility applications.
keywords:
[…]
@@ 1.1 Motivation
Human mobility modeling has emerged as a critical component in urban planning, transportation management, business, and public policy development (González et al., 2008; Liao et al., 2024; Stanford et al., 2024). Understanding how people move through cities and interact with the surrounding environment is fundamental to creating efficient, sustainable, and livable urban spaces. This importance has grown with increasing urbanization and the complexities of modern transportation systems (Hasan et al., 2013; Ma et al., 2024).
Foundation models and large language models (LLMs) offer transformative capabilities for mobility modeling by enabling multi-modal data integration, semantic enrichment, and cross-context generalization. These capabilities directly address the data fragmentation challenges that have historically limited mobility analysis, allowing for more comprehensive insights than previously possible with domain-specific approaches. Recent advances in multi-modal transportation systems have demonstrated the potential of sophisticated machine learning approaches, including graph neural networks for transportation recommendation systems (Yan et al., 2024) and ensemble learning methods for analyzing transportation behavior patterns (Yan et al., 2025).
One of the key research challenges in mobility modeling lies in integrating heterogeneous data sources to address incomplete or sparse datasets. Data fusion has emerged as a critical approach for combining diverse data types and enhancing modeling accuracy and robustness. Existing methods have demonstrated the potential of integrating sources such as GPS, social media, and environmental data to provide richer insights into mobility patterns (Ounoughi and Yahia, 2023; Afyouni et al., 2022). Researchers have also explored techniques for cross-domain integration and spatial-temporal modeling to improve movement predictions and scalability (Zheng, 2015; Wei et al., 2024). Although significant progress has been made, existing approaches still encounter challenges such as data sparsity, heterogeneity, and limited adaptability across diverse geographic contexts. To address these issues, this paper proposes a foundation model for mobility through cross-domain data fusion that enhances the applicability, precision, and generalizability of mobility modeling across diverse geographic and socio-demographic settings.
Figure 1: Integration of cross-domain multi-modal mobility data with human mobility modeling. (a) The need for data fusion and (b) the fused data.
@@ 1.2 Challenges
Comprehensive datasets addressing the diverse requirements of mobility research remain scarce. An ideal mobility dataset would integrate rich semantic information with high-resolution trajectory data and demographic attributes while ensuring privacy protection. However, as shown in Figure 1 (a), existing datasets often focus on only a subset of these dimensions (Kapp and Mihaljevic, 2023). The complexity of combining these diverse data sources stems from two key challenges: cross-domain data harmonization and multi-modal information integration, as shown in Figure 2.
Foundation models for mobility must address unique challenges in data integration that differ from those in natural language processing or computer vision. The multi-modal nature of mobility data—spanning continuous trajectories, discrete locations, semantic activities, and demographic attributes—requires sophisticated harmonization techniques to create coherent representations.
The fundamental challenge of cross-domain data fusion lies in combining data sources that complement each other but often contain conflicting information. Our foundation model addresses this by fusing data listed in Figure 2, where each source has distinct limitations: trajectories lack semantic context and are usually low quality, surveys are biased and expensive, and infrastructure data requires synthesis with human mobility patterns to be meaningful.
@@
To address these challenges, we propose a foundation model through cross-domain data fusion that synthesizes mobility data with integrated semantic, trajectory, demographic, and location-based information, as shown in Figure 1 (b). By leveraging the semantic understanding capabilities of LLMs and the generalization power of transfer learning, our approach enables a more comprehensive understanding of mobility patterns while preserving privacy.
@@ 1.3 Contribution
This research establishes a foundation model for human mobility that bridges multiple data gaps: between high-resolution but semantically sparse trajectories and context-rich but limited surveys, between population-level statistics and individual-level behaviors, and between static infrastructure data and dynamic human movements. Our approach enables new insights into human mobility and its implications for societal infrastructure, which hold promise for numerous applications, including traffic optimization, urban planning, and anomaly detection (He et al., 2024; Ma et al., 2024; Stanford et al., 2024).
Our research makes several significant contributions:
  • We developed a novel foundation model that processes and integrates cross-domain multi-modal data sources to create comprehensive synthetic datasets while balancing usability and privacy considerations. The framework harmonizes heterogeneous data types (spatial-temporal data: GPS (Global Positioning System) (Veraset, 2024) and POI (Point of Interest) (OpenStreetMap contributors, 2017), mobility data: Household Travel Survey (HTS) (Administration, 2017), demographic data: Time Use Survey (TUS) (U.S. BLS, 2024), and traffic data: road network data (OSM) (OpenStreetMap contributors, 2017)) and demonstrates generalization capabilities through validation with large-scale traffic simulation.
  • We incorporated large language models (LLMs) as a core component for semantic enrichment of trajectory data, fusing the realism of GPS data with the knowledge of survey data and reconstructing missing activities with a survey-trained model. This LLM-powered approach significantly improves the completeness and interpretability of mobility trajectories.
  • We implemented transfer learning mechanisms within our foundation model to address two key data challenges: 1) enabling knowledge and pattern adaptation in data-sparse regions through semi-supervised learning, leveraging limited GPS trajectories to adapt source domain patterns without requiring complete ground truth data, and 2) incorporating demographic variations into anonymized trajectories through unsupervised distribution adaptation, aligning mobility patterns with demographic characteristics without pairwise individual labels.
Figure 2: An overview of the proposed data fusion framework. The captions within the methodology indicate their corresponding sections in the paper for reference.
@@ 2.1 Human Travel Trajectory Reconstruction
Trajectory reconstruction has become crucial in understanding human mobility patterns, especially when dealing with incomplete datasets. Recent advancements have addressed the challenges of data sparsity and irregularity through innovative techniques. Chen et al. (2019) introduced the Context-enhanced Trajectory Reconstruction (CTR) method, using tensor factorization to reconstruct complete individual trajectories from sparse Call Detail Records. Li et al. (2019) proposed the Multi-criteria Data Partitioning Trajectory Reconstruction (MDP-TR) method for large-scale, low-frequency mobile phone datasets, enhancing reconstruction performance by considering spatiotemporal patterns of missing data and individual similarities. For GPS data, Zheng et al. (2010) developed a collaborative system for location and activity recommendations, demonstrating significant improvements in inferring activity types. Alexander et al. (2015) emphasized the importance of comprehensive temporal and spatial analysis in trajectory reconstruction using mobile phone data.
@@ 2.2 Large Language Models for Mobility Modeling
Large language models (LLMs) have demonstrated significant potential in mobility modeling by leveraging their ability to process contextual and semantic information from diverse data sources. Recent advancements, such as those by Liu et al. (2024) and Cai et al. (2024), highlight the versatility of LLMs in tasks like trajectory reconstruction and decision-making. Studies on foundational models like LLaMA (Touvron et al., 2023), PaLM (Anil et al., 2023), and TEG-DB (Li et al., 2024) further underline their power in multi-modal data fusion and semantic interpretation. These capabilities make LLMs valuable complements to traditional methods, particularly in addressing challenges of data sparsity and complex semantic understanding in mobility applications.
@@ 2.3 Transfer Learning
Howard and Ruder (2018) gradual unfreezing technique could be adapted to preserve general mobility patterns while adjusting to specific regional characteristics. Peters et al. (2019) and Merchant et al. (2020) suggest that middle layers are often the most transferable, which could be useful when adapting mobility models from data-rich to data-scarce regions. Liu et al. (2019) explored freezing bottom layers while fine-tuning top layers, a strategy that could preserve fundamental human movement patterns while adapting to unique characteristics of specific urban environments.
@@ 3 Problem Formulation
The transfer learning stage (4.2) is designed to capture real-world mobility dynamics while enabling deployment in data-scarce regions. Through knowledge (4.2.1) and pattern (4.2.2) adaptation, the model can be adapted to different regions and cultures and learns to balance the generic knowledge learned from surveys with the realism in GPS trajectories, including region-specific patterns. Our demographic adaptation module further enhances realism by incorporating sociodemographic variations using time use survey (TUS) data, which provides rich activity patterns across different population groups, ensuring our synthetic trajectories reflect both demographic-specific behaviors and real-world mobility dynamics while maintaining privacy.
Finally, the third stage (4.3) grounds these behaviorally realistic patterns in physical space by fusing the generated activity chains with spatial distributions and traffic data to assign feasible locations, enabling validation through large-scale traffic simulation. This complete integration of dynamic human behavior with static infrastructure data creates a comprehensive framework that generates mobility patterns reflecting both the complexity of human decision-making and the constraints of physical infrastructure.
@@
With our semantically annotated trajectories represented as $A_{j}$, we propose an activity reconstruction model $M^{R1}$ tailored to a specific region $R1$. For an incomplete $A_{j}$, the model synthesizes a new activity chain $A_{j}^{\prime}=M^{R1}(A_{j})$, by filling the gaps based on the common activity patterns learned from the region $R1$.
The activity reconstruction adapts cultural effects to activity patterns through transfer learning, but incorporating sociodemographic correlation patterns poses a challenge. To address this, we introduce a demographic adaptation module that modifies reconstructed activity chains based on sociodemographic characteristics. Let us define the sociodemographic space $\mathcal{S}$ as: $\mathcal{S}=\{\mathbf{s}_{g}|g\in\mathcal{G}\}$, where $\mathcal{G}$ represents the set of demographic groups. For a demographic space with $N$ dimensions (e.g., age, gender, employment status), where each dimension $i$ has $C_{i}$ categories, the total number of demographic groups is: $|\mathcal{G}|=\prod_{i=1}^{N}C_{i}$.
The adapter functions as an additional layer after the reconstruction model $M^{R1}$. Given a batch of reconstructed activity chains probability matrix $\mathbf{Y}\in\mathbb{R}^{B\times K\times|T|}$, where $B$ is the batch size, $K=96$ represents 96 time steps with 15-minute time intervals for a day, and $|T|$ represents the number of activity types defined in Table 1, we introduce an adapter matrix $\mathbf{M}\in\mathbb{R}^{|\mathcal{G}|\times|T|}$. For each demographic group $g\in\mathcal{G}$, its adapter vector $\mathbf{a}_{g}\in\mathbb{R}^{|T|}$ modifies the reconstructed activity probabilities. The demographically-adapted activity probabilities for group $g$ across $K$ time intervals are computed as:
$\mathbf{P}_{g}=\text{softmax}(\mathbf{Y}+\mathbf{a}_{g},\text{axis}=|T|)$ (1)
where $\mathbf{a}_{g}$ is broadcast across the $K$ time intervals.
Finally, we assign the activity chains to physical zonal locations while preserving spatial and temporal feasibility. Given a set of TAZs $\mathcal{Z}$ and road network $\mathcal{R}$, we assign locations to activities by optimizing:
$L^{*}=\operatorname*{arg\,min}_{L\in\mathcal{Z}}\sum_{i=1}^{n}\sum_{j=1}^{m_{i}}d(L_{i,j},L_{i,j+1})$ (2)
subject to constraints on travel time feasibility via $\mathcal{R}$ and consistency with observed spatial distributions. This enables validation against real-world traffic patterns.
@@ 4.1 Generative Data Fusion
4.1.1 LLM-Informed Trajectory Semantic Annotation
Human trajectory data collected from GPS devices include geo-location and time-stamped stay points that elucidate individual mobility patterns. However, these trajectories suffer from both semantic sparsity and temporal incompleteness. We develop the generation-based data fusion module, as Figure 2 (4.1), to address these two challenges.
Adopting the methodology from our previous work  (Liu et al., 2024), we use LLMs to annotate stay points using the POI dataset. This process utilizes LLM to assign semantic labels to POIs based on attributes such as "name", "amenity", and "construction type". Each POI is then categorized into one of 15 activity codes as described in Table 1. The LLM’s prompt configuration includes a "Task Description", "Activity Code Illustration", and a description of each POI in natural language format, ensuring comprehensive understanding and accurate classification. The output from the LLM provides the top three most relevant activity codes for each POI, each with a corresponding probability.
@@ 4.1.1 LLM-Informed Trajectory Semantic Annotation
4.1.2 Data Augmentation by Activity Reconstruction
While this LLM-based annotation provides semantic context to the trajectories, their inherent incompleteness limits their direct use for mobility analysis. Therefore, our framework continues with an activity reconstruction process that fuses these annotated but incomplete trajectories with comprehensive mobility data from travel surveys. A base Activity Reconstruction Model $M^{0}$ is trained on survey data to learn generic activity patterns and generate complete activity chains based on limited observed GPS stay points. It can be adapted to other regions with a small dataset, supported by the strong similarity in human activities across regions  (Schneider et al., 2013; Cao et al., 2019). The adaptation of this base model to realistic geographical and cultural contexts is detailed in Section 4.2.